# Sharp transitions in small exciton spectra for multi-orbital lattice systems

### v2

@@ VI Acknowledgments
We acknowledge support from the Natural Sciences and Engineering Research Council of Canada (NSERC). We gratefully acknowledge the use of computing resources from the Stewart Blusson Quantum Matter Institute computing cluster LISA. Lastly, we also acknowledge that the research is done on the UBC Point Grey (Vancouver) campus, which sits on the traditional, ancestral, unceded territory of the $x^{w}m$$\theta k^{w}$$\overset{,}{y}$@$m$ (Musqueam) First Nation.
@@ Appendix A Hoppings in the 2D, three-orbital valence band
For the specific calculations shown in the main paper, we set $t_{v3}=1$, the electron hopping in the conduction band is $t_{c}=1$, and $\Delta=16.5$.
@@ Appendix B Recurrence relations and their convergence
$\bm{G}_{\gamma,\delta}(K,z)\equiv\begin{bmatrix}G_{\gamma,s}(K,\delta,z)\\[2.0pt] G_{\gamma,p}(K,\delta,z)\end{bmatrix}$.
We start by considering the action of the Hamiltonian $H$ on a ket. For $\delta\neq 0$, the on-site part of $H$ leaves $\delta$ unchanged, while hopping changes $\delta\to\delta\pm 1$.
@@
These are solved in terms of the ansatz $\bm{G}_{\gamma,\delta}(K,z)=\underline{S}_{\pm}\bm{G}_{\gamma,\delta\mp 1}(K,z)$, where the upper/lower signs correspond to $\delta>0$ and $\delta<0$, respectively, and $\underline{S}_{+}=[\underline{\alpha}-\underline{\beta}\underline{S}_{+}]^{-1}\underline{\gamma}$ and $\underline{S}_{-}=[\underline{\alpha}-\underline{\gamma}\underline{S}_{-}]^{-1}\underline{\beta}$ are continued fractions of matrices, calculated with $M$ iterations starting from ${\underline{S}}_{\pm}=0$. The cutoff $M$ is chosen sufficiently large to achieve convergence to the desired accuracy (see below). Physically, $M$ defines a variational solution whereby the electron and hole are not allowed to be $M$ or more sites apart.
@@
The method described above can be generalized straightforwardly to lattices in higher dimensions [35]. Here we briefly describe the approach for a triangular lattice. The relative distance between electron and hole is $\bm{\delta}\in\mathbb{Z}^{2}$ and we define $N=|\bm{\delta}|$ (in units of $a$) as the minimum number of hops to get from an onsite configuration ($\delta=0$) to the relative separation $\bm{\delta}$. For a triangular lattice there are $6N$ distinct $\bm{\delta}$} separations for a given $N$, lying on a circular shell. We index these counter-clockwise starting from the one lying along the +x+x-axis, and group together into $\bm{G}_{\gamma,N}$ the $d_{N}=18N$ Green's functions with the same $N$. The extra factor of 3 is because we consider 3 possible valence orbitals per site.
@@
The matrices $\underline{\alpha}_{N}\in\mathbb{C}^{3d_{N}\times 3d_{N}}$, $\underline{\beta}_{N}\in\mathbb{C}^{3d_{N}\times 3d_{N+1}}$, $\underline{\gamma}_{N}\in\mathbb{C}^{3d_{N}\times 3d_{N-1}}$ can be read off from Eq. (13). Eq. (24) is solved using the ansatz $\bm{G}_{\gamma,N}={\underline{S}}_{N}(\mathbf{K},z)\bm{G}_{\gamma,N-1}$, where ${\underline{S}}_{N}(\mathbf{K},z)=[\underline{\alpha}_{N}(\mathbf{K},z)-\underline{\beta}_{N}(\mathbf{K}){\underline{S}}_{N+1}(\mathbf{K},z)]^{-1}\underline{\gamma}_{N}(\mathbf{K})$ is calculated iteratively starting from $\underline{S}_{M}(\mathbf{K},z)=0$ for a sufficiently large $M$ such that the desired accuracy is reached (see below). In particular, this allows us to calculate $\underline{S}_{1}(\mathbf{K},z)$.
@@
Figure 9: Relative error $\epsilon_{M}=|E_{\rm exc}(M)/E_{\rm final}-1|$ as a function $M$, where $E_{\text{exc}}(M)$ is the lowest exciton energy calculated using truncation at $M$, while $E_{\text{final}}$ is the converged value for the lowest exciton energy, obtained by truncating at a sufficiently large cutoff such that increasing it by 1 results in a relative change below $10^{-6}$. (a) 1D results: blue circles correspond to a large $U_{p}=U_{s}=2.2$, while red triangles correspond to a weaker $U_{p}=U_{s}=0.8$. In both cases $t_{c}=1,t_{ss}=1.1,t_{sp}=0.5,t_{pp}=-1.0$. The former case has a smaller exciton (the dashed vertical lines show the corresponding exciton radius $\xi$) and thus converges faster with increasing $M$; (b) corresponding 2D results for the three band model with $U_{x^{2}-y^{2}}=U_{xy}=U_{3z^{2}-r^{2}}=12$ and the other parameters listed in Appendix A. Here $\xi\approx 0.856$ (not shown) is very small and the convergence with M is very rapid.
@@
Truncating at shell radius $N=1$ (i.e. setting $M=2$ when calculating the continued fractions $\underline{S}_{N}$), gives the one-shell approximation used in Appendix G. The cut-off used for Fig. 4-8 in the main text is $M=10$.
The convergence in terms of the cutoff $M$ is illustrated in Fig. 9 for sample 1D (top) and 2D (bottom) cases. The stronger the attraction and therefore the smaller the exciton, the faster the convergence with increasing $M$. For really small excitons, even a cutoff $M=2$ can already give reasonable accuracy, as further discussed in Appendix G. This illustrates the main advantage of using this method for strongly bound excitons.
@@ Appendix C Longer-range Coulomb attraction
Figure 10: Heat map similar to Fig. 4 showing the spectral weight $A_{\alpha\alpha}({\bf K},\omega)=-{1\over\pi}\text{Im}G_{\alpha\alpha}({\bf K},\bm{0},z)$ for $\bf K$} along high-symmetry lines in the Brillouin zone, summing over all the source orbitals. Parameters are: on-site attraction $U=11$, nearest-neighbor attraction $U_{1}=2$, all other parameters are as defined in Appendix A.
As an example, in Fig. 10 we show the exciton spectrum when a nearest neighbor attraction $U_{1}=2$ is included beside the on-site $U=11$; all other parameters are as listed in Appendix A. Importantly, we find that adding a nearest-neighbor attraction preserves the sharp ground-state momentum transition discussed in the main text, and in fact it provides additional tunability to achieve it. Arbitrarily long-range attraction can be added and studied similarly, however very small excitons only 'sample' short relative distances so whether there is attraction at distances larger than the cutoff $M$ is irrelevant.
@@
Figure  12 shows a comparison between the lowest-energy exciton binding energy predicted by the 2D lattice model (red circles) and by its corresponding continuum model (blue solid line). The parameters for the lattice model are $t_{c}=1,t_{v1}=0,t_{v2}=0,t_{v3}=1,t_{v4}=0,\Delta=7$, but qualitatively similar results are obtained for any other choice of parameters. For the continuum model, we use the lattice results for weak attraction $U\in[0,2]$ to fit $R_{0}$ and $g$ from Eq. (32), and then generate the continuum prediction for all $U$.
Figure 12: Comparison between the lowest-exciton binding energy predicted by the 2D lattice model (red circles) and by the continuum model (blue line). Here, the parameters for the lattice model are $t_{c}=1,t_{v1}=0,t_{v2}=0,t_{v3}=1,t_{v4}=0,\Delta=7$, while those for the continuum model are obtained by fitting the lattice results with Eq. (32) for $U\in[0,2]$. The comparison is plotted (a) on a semi-log scale vs. $1/U$; and (b) on linear scale vs. $U$. As expected, the agreement is excellent for weak attraction $U$ but significant deviations appear for stronger attraction $U$. See text for more details.
@@ Appendix G Minimal variational approximation for the 2D strongly bound excitons
If the Coulomb attraction is so strong that the exciton is Frenkel-like, we expect a reasonable estimate to be obtained if we restrict the relative distance between electron and hole to be $|{\bm{\delta}}|\leq a$. One could then diagonalize the full Hamiltonian within this variational space, but the same results can be obtained efficiently from our method by setting all the propagators $\bm{G}_{\gamma,N}=0$ for $N\geq 2$, i.e. by choosing the cutoff at $M=2$.

### v3

@@ VI Acknowledgments
This project was undertaken thanks in part to funding from the Max Planck-UBC-UTokyo Center for Quantum Materials and the Canada First Research Excellence Fund, Quantum Materials and Future Technologies Program, as well as the Natural Sciences and Engineering Research Council of Canada (NSERC). We gratefully acknowledge the use of computing resources from the Stewart Blusson Quantum Matter Institute computing cluster LISA. Lastly, we also acknowledge that the research is done on the UBC Point Grey (Vancouver) campus, located on the traditional, ancestral, unceded territory of the $x^{w}m$$\theta k^{w}$$\overset{,}{y}$@$m$ (Musqueam) First Nation.
@@ Appendix A Hoppings in the 2D, three-orbital valence band
For the specific calculations shown in the main paper, we set $t_{v3}=1$, the electron hopping in the conduction band is $t_{c}=1$, and $\Delta=15.5$.
@@ Appendix B Recurrence relations and their convergence
$\bm{G}_{\gamma,\delta}(K,z)\equiv\begin{bmatrix}G_{\gamma,s}(K,\delta,z)\\[2.0pt] G_{\gamma,p}(K,\delta,z)\end{bmatrix}$. (19)
We start by operating the Hamiltonian $H$ on a basis state defined in Eq. 7. For $\delta\neq 0$, the on-site part of $H$ leaves $\delta$ unchanged, while hopping changes $\delta\to\delta\pm 1$.
@@
These are solved in terms of the ansatz $\bm{G}_{\gamma,\delta}(K,z)=\underline{S}_{\pm}\bm{G}_{\gamma,\delta\mp 1}(K,z)$, where the upper/lower signs correspond to $\delta>0$ and $\delta<0$, respectively, and $\underline{S}_{+}=[\underline{\alpha}-\underline{\beta}\underline{S}_{+}]^{-1}\underline{\gamma}$ and $\underline{S}_{-}=[\underline{\alpha}-\underline{\gamma}\underline{S}_{-}]^{-1}\underline{\beta}$ are continued fractions of matrices, calculated with $M$ iterations starting from ${\underline{S}}_{\pm}=0$. The cutoff $M$ is chosen sufficiently large to achieve convergence to the desired accuracy (see below). Physically, $M$ defines a variational solution whereby the electron and hole are not allowed to be $M$ or more sites apart, i.e. setting to zero all propagators with $|\delta|\geq M$. We note that in the main text, we called this cut-off $\delta_{M}$ instead of $M$.
@@
The method described above can be generalized straightforwardly to lattices in higher dimensions [48]. Here we briefly describe the approach for a triangular lattice. The relative displacement between electron and hole is $\bm{\delta}\in\mathbb{Z}^{2}$ and we define $N=|\bm{\delta}|$ (in units of $a$) as the minimum number of hops to get from an onsite configuration ($\delta=0$) to the relative separation $\bm{\delta}$. For a triangular lattice there are $6N$ distinct $\bm{\delta}$} separations for a given $N$, lying on a hexagonal shell. We index these counter-clockwise starting from the one lying along the +x+x-axis, and group together into $\bm{G}_{\gamma,N}$ the $d_{N}=18N$ Green's functions with the same $N$. The extra factor of 3 is because in our model we consider 3 valence orbitals and 1 conduction orbital per site; generalization to other numbers of valence and/or conduction orbitals is straightforward.
@@
The matrices $\underline{\alpha}_{N}\in\mathbb{C}^{d_{N}\times d_{N}}$, $\underline{\beta}_{N}\in\mathbb{C}^{d_{N}\times d_{N+1}}$, $\underline{\gamma}_{N}\in\mathbb{C}^{d_{N}\times d_{N-1}}$ can be read off from Eq. (15). Eq. (27) is solved using the ansatz $\bm{G}_{\gamma,N}={\underline{S}}_{N}(\mathbf{K},z)\bm{G}_{\gamma,N-1}$, where ${\underline{S}}_{N}(\mathbf{K},z)=[\underline{\alpha}_{N}(\mathbf{K},z)-\underline{\beta}_{N}(\mathbf{K}){\underline{S}}_{N+1}(\mathbf{K},z)]^{-1}\underline{\gamma}_{N}(\mathbf{K})$ is calculated iteratively starting from $\underline{S}_{M}(\mathbf{K},z)=0$ for a sufficiently large cutoff $M$ such that the desired accuracy is reached (see example below). In particular, this allows us to calculate $\underline{S}_{1}(\mathbf{K},z)$. For clarity, we note that $M$ is related to the generic cutoff $\delta_{M}$ used in the main text; the explicit relation depends on the lattice symmetry.
@@
Figure 10: Relative error $\epsilon_{M}=|E_{\rm exc}(M)/E_{\rm final}-1|$ as a function of $M$, where $E_{\text{exc}}(M)$ is the lowest exciton energy calculated using truncation at $M$, while $E_{\text{final}}$ is the converged value for the lowest exciton energy, obtained by truncating at a sufficiently large cutoff $M$ such that increasing it by 1 results in a relative change below $10^{-6}$. (a) 1D results: blue circles correspond to a large $U_{p}=U_{s}=2.2$, while red triangles correspond to a weaker $U_{p}=U_{s}=0.8$. In both cases $t_{c}=1,t_{ss}=1.1,t_{sp}=0.5,t_{pp}=-1.0$. The former case has a smaller exciton (the dashed vertical lines show the corresponding exciton radius $\xi$) and thus converges faster with increasing $M$; (b) corresponding 2D results for the three band model with $U_{x^{2}-y^{2}}=U_{xy}=U_{3z^{2}-r^{2}}=12$ and the other parameters listed in Appendix A. Here $\xi\approx 0.856$ (not shown) is very small and the convergence with $M$ is very rapid.
@@
Truncating at shell radius $N=1$ (i.e. setting $M=2$ when calculating the continued fractions $\underline{S}_{N}$), gives the one-shell approximation used in Appendix G. The cut-off used for Fig. 5-9 in the main text is $M=10$.
The convergence in terms of the cutoff $M$ is illustrated in Fig. 10 for sample 1D (top) and 2D (bottom) cases. The stronger the attraction and therefore the smaller the exciton, the faster the convergence with increasing $M$. For really small, Frenkel-like excitons, even a cutoff $M=2$ can already give reasonable accuracy, as further discussed in Appendix G. This illustrates the main advantage of using this method for strongly bound excitons.
@@ Appendix C Longer-range Coulomb attraction
Figure 11: Heat map similar to Fig. 4 showing the spectral weight $A_{\alpha\alpha}({\bf K},\omega)=-\frac{1}{\pi}\text{Im}G_{\alpha\alpha}({\bf K},\bm{0},z)$ for $\bf K$} along high-symmetry lines in the Brillouin zone, summing over all the source orbitals. Parameters are: on-site attraction $U=11$, nearest-neighbor attraction $U_{1}=2$, all other parameters are as defined in Appendix A.
As an example, in Fig. 11 we show the exciton spectrum when a nearest neighbor attraction $U_{1}=2$ is included beside the on-site $U=11$; all other parameters are as listed in Appendix A. Importantly, we find that adding a nearest-neighbor attraction preserves the sharp ground-state momentum transition discussed in the main text, and in fact it provides additional tunability to achieve it. Arbitrarily long-range attraction can be added and studied similarly. However, very small excitons only "sample" short relative displacements, so whether there is attraction at displacements larger than the cutoff $M$ is irrelevant.
@@
Figure  1314 show comparisons between the lowest-energy exciton binding energy predicted by the 2D lattice model (red circles) and by its corresponding continuum model (blue solid line). The parameters for the lattice model are $t_{c}=1,t_{v1}=0,t_{v2}=0,t_{v3}=1,t_{v4}=0,\Delta=7$, but qualitatively similar results are obtained for any other choice of parameters. For the continuum model, we use the lattice results for weak attraction $U\in[0,2]$ to fit $R_{0}$ and $g$ from Eq. (35), and then generate the continuum prediction for all $U$.
Figure 13: Comparison between the lowest-exciton binding energy predicted by the 2D lattice model (red circles) and by the continuum model (blue line). Here, the parameters for the lattice model are $t_{c}=1,t_{v1}=0,t_{v2}=0,t_{v3}=1,t_{v4}=0,\Delta=7$, while those for the continuum model are obtained by fitting the lattice results with Eq. (35) for $U\in[0,2]$. The comparison is plotted on a semi-log scale vs. $1/U$. See text for more details.
Figure 14: Same comparison as in Fig. 13 but on linear scale vs. $U$. As expected, the agreement is excellent for weak attraction $U$ but significant deviations appear for stronger attraction $U$.
@@ Appendix G Minimal variational approximation for the 2D strongly bound excitons
If the Coulomb attraction is so strong that the exciton is Frenkel-like, we expect a reasonable estimate to be obtained if we restrict the relative displacement between electron and hole to be $|{\bm{\delta}}|\leq a$. One could then diagonalize the full Hamiltonian within this variational space, but the same results can be obtained efficiently from our method by setting all the propagators $\bm{G}_{\gamma,N}=0$ for $N\geq 2$, i.e. by choosing the cutoff at $M=2$.